# Structured dataset of reported cloud seeding activities in the United States (2000–2025) using a large language model

### v2

@@ Abstract
Cloud seeding, a weather modification technique aimed at locally enhancing precipitation, has been used in the Western United States since the 1940’s. However, datasets to facilitate large-scale analysis of cloud seeding efforts are not currently available. To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S. from 2000 to 2025, including the year, season, state, seeding agent, apparatus used for deployment, and purpose. Using OpenAI’s o4-mini large language model (LLM), combined with multi-stage PDF-to-text conversion and response-parsing code, we processed 836 historical reports from the National Oceanic and Atmospheric Administration (NOAA) to extract the data. The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo. Our results address the gap in accessible cloud seeding data in the United States, and demonstrate the potential for LLMs to extract structured environmental data from historical documents. More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
@@
Cloud seeding is a weather modification technique used to enhance precipitation, typically in regions affected by drought, water scarcity, or to support snowpack accumulation.
In the United States, cloud seeding has been practiced since the 1940s but reporting was not standardized until the Weather Modification Reporting Act of 1972 [1]. The Act requires individuals and organizations to notify the U.S. Department of Commerce at least 10 days before and after conducting weather modification operations. These reports are submitted to NOAA via email and compiled on the NOAA Weather Modification website [2].
@@ Methods
The final cloud seeding dataset is provided as a single CSV file titled cloud_seeding_us_2000_2025.csv, using one row per cloud seeding project and one column per extracted field. It is publicly available on Zenodo at the following DOI: https://doi.org/10.5281/zenodo.14925812. It contains structured information extracted from 836 cloud seeding project reports in the United States between 2000 and 2025.
@@
To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM. The review was manually conducted independently by two human experts. A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level. For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-golden.py.
@@
We evaluated several OpenAI models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3 [10]. This comparison allowed us to assess how accuracy varied with model complexity and specialization. Among these, o4-mini offered the best balance of performance and cost, achieving 95.00% accuracy at approximately \$0.005 per document. While o3 achieved the highest accuracy at 96.33%, it was an order of magnitude more expensive, costing around \$0.05 per document.
@@
We also evaluated how different prompts influenced extraction accuracy, holding the model constant (o4-mini). Three prompt formats were tested: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and a chain-of-thought format (Prompt C) that encouraged step-by-step reasoning. Prompt C achieved the highest accuracy at 95.00%, suggesting that explicit reasoning steps can substantially improve structured data extraction.
@@
Together, these validation efforts demonstrate that LLM-based data extraction can be significantly improved with thoughtful prompt engineering. Chain-of-thought reasoning and example-driven prompts, in particular, are well-suited for scientific data extraction from complex archival documents. While our evaluation covered key variants, further experimentation could uncover additional insights.

### v3

@@ Abstract
Cloud seeding, a weather modification technique used to increase precipitation, has been employed in the western United States since the 1940s. However, structured datasets are not currently available to analyze these efforts. To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S. from 2000 to 2025, including the year, season, state, seeding agent, apparatus used for deployment, and purpose. Using OpenAI’s o4-mini large language model (LLM), combined with multi-stage PDF-to-text conversion and response-parsing code, we processed 836 historical reports from the National Oceanic and Atmospheric Administration (NOAA) to extract the data. The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo. Our results help fill the gap in structured cloud seeding data and demonstrate the potential for LLMs to extract structured environmental data from historical documents. More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
@@
Cloud seeding is a weather modification technique used to enhance precipitation, typically in regions affected by drought, water scarcity, or to support snowpack accumulation. In the United States, cloud seeding has been practiced since the 1940s but reporting was not standardized until the Weather Modification Reporting Act of 1972 [1]. The Act requires individuals and organizations to notify the U.S. Department of Commerce at least 10 days before and after conducting weather modification operations. These reports are submitted to NOAA via email and compiled on the NOAA Weather Modification website [2].
@@ Methods
The final cloud seeding dataset is provided as a single CSV file titled cloud_seeding_us_2000_2025.csv, using one row per cloud seeding project and one column per extracted field (Figure 2). It is publicly available on Zenodo at the following DOI: https://doi.org/10.5281/zenodo.14925812. It contains structured information extracted from 836 cloud seeding project reports in the United States between 2000 and 2025.
@@
To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM (Table 1). The review was manually conducted independently by two human experts. A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level. For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-golden.py [9].
@@
We evaluated several OpenAI models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3 [10] (Table 2). This comparison allowed us to assess how accuracy varied with model complexity and specialization. Among these, o4-mini offered the best balance of performance and cost, achieving 95.00% accuracy at approximately \$0.005 per document. While o3 achieved the highest accuracy at 96.33%, it was an order of magnitude more expensive, costing around \$0.05 per document.
@@
We also evaluated how different prompts influenced extraction accuracy, holding the model constant with o4-mini (Table 3). Three prompt formats were tested: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and a chain-of-thought format (Prompt C) that encouraged step-by-step reasoning. Prompt C achieved the highest accuracy at 95.00%, suggesting that explicit reasoning steps can substantially improve structured data extraction.
@@
Together, these validation efforts demonstrate that LLM-based data extraction can be significantly improved with thoughtful prompt engineering. In particular, chain-of-thought reasoning and example-driven prompts are well-suited for scientific data extraction from complex historical documents.